OpenStudy (anonymous):

x+y=6 and x+4y=3 solve by substitution

OpenStudy (anonymous):

x=6-y x+4y=3 6-y+4y=3 3y=-3 y=-1 x=7

OpenStudy (anonymous):

x=7 y=-1